OpenStudy (anonymous):

Please help 3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. -6y = -12 -6y = 12 -2y = -12 2y = 12

OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 -->(-1) ------------- 3x - 4y = 7 -3x - 2y = 5 (result of multiplying by -1) ------------ -6y = 12